# VLMLight: Traffic Signal Control via Vision-Language Meta-Control and Dual-Branch Reasoning

### v1

@@ 1 Introduction
The emergence of large language models (LLMs) achiam2023gpt has opened new possibilities for incorporating high-level reasoning and generalization into traffic signal control systems tang2024large ; lai2025llmlight . Recent efforts have explored the use of LLMs to handle edge cases and long-tail scenarios that are difficult for RL-based policies to address wang2024llm ; pang2024illm . However, current LLM-based approaches often rely on templated or manually crafted textual descriptions of traffic scenes, which lack the richness and fidelity needed to capture complex, real-world dynamics, leading to significant information loss, especially in visually grounded or ambiguous situations. Inference speed poses another hurdle: multi-step reasoning typically makes LLMs orders of magnitude slower than RL controllers, rendering LLMs impractical as standalone controllers in traffic environments where both latency and precision are essential. Moreover, current TSC simulators, such as the widely used SUMO SUMO2018 and CityFlow tang2019cityflow , can only provide statistical data for traffic intersections and cannot render real-time images, limiting the possibility to fuse the visual domain with language-based reasoning for traffic signal control.
@@ 3.1 Scene Understanding via VLM
[…]
@@ 4.3 Ablation Study
Table 2: Ablation results on Songdo, impact of structured reasoning components and corresponding inference time..
@@ 5 Related Work
TSC methods fall into three categories: rule-based, RL-based, and LLM-based. Traditional approaches like Webster’s method webster1958traffic , SOTL cools2008self , and MaxPressure varaiya2013max use fixed heuristics and perform well under steady traffic, but struggle with real-time adaptability wei2019survey . RL-based methods aslani2017adaptive ; chu2019multi ; wang2024unitsa ; chen2023efficient ; du2023safelight ; gu2024pi ; chen2024difflight ; yang2024mallight improve responsiveness by learning from interaction, but often overlook safety-critical events due to simplified states and static rewards yau2017survey ; qadri2020state ; wei2021recent ; zhao2024survey . Recent LLM-based approaches tang2024large ; wang2024llm ; lai2025llmlight introduce high-level reasoning for rare or long-tail cases, but rely on hand-crafted text inputs and lack true visual grounding. Moreover, they underperform in routine traffic control compared to specialized RL agents.
@@
Most TSC simulators, such as SUMO SUMO2018 , CityFlow tang2019cityflow , and LibSignal mei2024libsignal , provide only structured traffic states (e.g., vehicle counts, signal phases) without visual outputs, limiting their use in vision-grounded reasoning. High-fidelity driving simulators such as CARLA Dosovitskiy17 and MetaDrive li2022metadrive offer photorealistic rendering, yet their design focuses on autonomous vehicle tasks rather than intersection control. These platforms often lack native support for signal scheduling and impose significant overhead in scene customization and control integration.
@@ A.1 Details of Routine Control Policy
where each row $\mathbf{Z}_{t,k}\in\mathbb{R}^{d}$ corresponds to the enriched representation of time step $t-k$. To obtain a fixed-length state embedding for decision-making, we apply average pooling over the temporal axis:
@@ A.2 Algorithm for VLMLight
[…]
@@ B.1 Experiment Settings
For the RL-based components of VLMLight, we use the PPO [27] algorithm, implemented via the Stable Baselines3 library. To speed up training and improve the exploration of different traffic conditions, we deploy $30$ parallel processes, each interacting with a separate instance of the simulation. The total number of environment steps is set to $53e$ 5 and the batch size is configured to $64$. The learning rate follows a linear schedule, starting at $31e-$ 3 and gradually decreasing as the number of training steps increases. Additionally, the trajectory memory size is set to $3000$.

### v2

# VLMLight: Safety-Critical Traffic Signal Control via Vision-Language Meta-Control and Dual-Branch Reasoning Architecture

@@ 1 Introduction
The emergence of large language models (LLMs) achiam2023gpt has opened new possibilities for incorporating high-level reasoning and generalization into traffic signal control systems tang2024large ; lai2025llmlight . Recent efforts have explored the use of LLMs to handle edge cases and long-tail scenarios that are difficult for RL-based policies to address wang2024llm ; pang2024illm ; zou2025traffic . However, current LLM-based approaches often rely on templated or manually crafted textual descriptions of traffic scenes, which lack the richness and fidelity needed to capture complex, real-world dynamics, leading to significant information loss, especially in visually grounded or ambiguous situations. Inference speed poses another hurdle: multi-step reasoning typically makes LLMs orders of magnitude slower than RL controllers, rendering LLMs impractical as standalone controllers in traffic environments where both latency and precision are essential. Moreover, current TSC simulators, such as the widely used SUMO SUMO2018 and CityFlow tang2019cityflow , can only provide statistical data for traffic intersections and cannot render real-time images, limiting the possibility to fuse the visual domain with language-based reasoning for traffic signal control.
@@ 3.1 Scene Understanding via VLM
$\{T_{1},T_{2},\dots,T_{D}\}={\texttt{Agent}\textsubscript{Scene}}(\{I_{1},I_{2},\dots,I_{D}\}),$ (1)
@@ 4.3 Ablation Study
Table 2: Ablation results on Songdo, impact of structured reasoning components and corresponding inference time.
@@ 5 Related Work
TSC methods fall into three categories: rule-based, RL-based, and LLM-based. Traditional approaches like Webster’s method webster1958traffic , SOTL cools2008self , and MaxPressure varaiya2013max use fixed heuristics and perform well under steady traffic, but struggle with real-time adaptability wei2019survey . RL-based methods aslani2017adaptive ; chu2019multi ; wang2024unitsa ; chen2023efficient ; du2023safelight ; gu2024pi ; chen2024difflight ; yang2024mallight improve responsiveness by learning from interaction, but often overlook safety-critical events due to simplified states and static rewards yau2017survey ; qadri2020state ; wei2021recent ; zhao2024survey . Recent LLM-based approaches tang2024large ; wang2024llm ; lai2025llmlight ; zou2025traffic introduce high-level reasoning for rare or long-tail cases, but rely on hand-crafted text inputs and lack true visual grounding. Moreover, they underperform in routine traffic control compared to specialized RL agents.
@@
Most TSC simulators, such as SUMO SUMO2018 , CityFlow tang2019cityflow , and LibSignal mei2024libsignal , provide only structured traffic states (e.g., vehicle counts, signal phases) without visual outputs, limiting their use in vision-grounded reasoning. High-fidelity driving simulators such as CARLA Dosovitskiy17 and MetaDrive li2022metadrive offer photorealistic rendering, but they are primarily designed for autonomous driving tasks rather than intersection control, and they often lack native support for signal scheduling or require extensive customization. To address this gap, SynTraC chen2024syntrac introduces an image-based dataset for TSC built upon CARLA, providing intersection images annotated with traffic states, signal phases, and reward information under diverse weather and lighting conditions. Although SynTraC represents a significant step toward visual perception in TSC, it remains limited to single-intersection scenarios and does not support user-defined or multi-view configurations, constraining its scalability and applicability to broader traffic management research.
@@ A.1 Details of Routine Control Policy
where each row $\mathbf{Z}_{t,k}\in\mathbb{R}^{d}$ corresponds to the enriched representation of time step tkt-k. To obtain a fixed-length state embedding for decision-making, we apply average pooling over the temporal axis:
@@ A.2 Algorithm for VLMLight
Algorithm 1 outlines the inference procedure of VLMLight. At each decision interval, the system receives multi-view images from the simulator and invokes the AgentScene to generate directional scene descriptions. A safety-prioritized meta-controller AgentModeSelector then determines whether to proceed with the fast RL policy or activate the structured reasoning branch. In routine conditions, a lightweight RL agent issues a control action based on the current traffic state. In contrast, for safety-critical scenarios, three LLM agents collaborate sequentially: the AgentPhase module transforms scene descriptions into phase-level summaries using a predefined phase-to-lane mapping $\mathcal{M}{\text{ph-lane}}$; the AgentPlan agent proposes a candidate action aligned with system objectives; and the AgentCheck agent verifies the action’s legality against the current feasible phase set $\mathcal{A}_{t}$. If the verification fails after $N_{\text{check}}$ attempts, the system falls back to the action proposed by AgentRL to ensure continued operation. This strict validation pipeline, coupled with the fallback mechanism, safeguards against potential noise or reasoning errors, ensuring that only feasible and safety-compliant actions are executed. The selected action is then executed in the simulator, and the simulation clock advances by a fixed interval $\Delta t$. This loop continues until the maximum simulation time $T{\max}$ is reached.
@@ B.1 Experiment Settings
For the RL-based components of VLMLight, we use the PPO [28] algorithm, implemented via the Stable Baselines3 library. To speed up training and improve the exploration of different traffic conditions, we deploy $30$ parallel processes, each interacting with a separate instance of the simulation. The total number of environment steps is set to $3e5$ and the batch size is configured to $64$. The learning rate follows a linear schedule, starting at 1e31e-3 and gradually decreasing as the number of training steps increases. Additionally, the trajectory memory size is set to $3000$.